OpenStudy (anonymous):

A sector has an area of 40cmsquared, and an angle of 30 degrees. The arc length of the sector, to 2 decimal points is: a)1.64cm b)2.66cm c)4.83cm d)6.47cm e)12.36cm

OpenStudy (anonymous):

\[r^2/2*\theta \] is the area of a sector formula

OpenStudy (anonymous):

arc length formula is \[2\pi*r*\theta\]

Directrix (directrix):

d)6.47cm 30/360 = 40(/pi r^2) --> this gives the portion or the circle's area covered by the sector 1/12 = 40/(3.14 r^2) 3.14 r^2 = 12*40 r^2 = 480/3.14 r ^2 = 152.866 r = 12.36 -------------- 30/360 = AL/[2 pi (12.36)] --> this gives the portion of the circumference length covered by the arc where AL represents the length of the arc 1/12 = AL/77.645 AL = 6.470

OpenStudy (anonymous):

@Directrix: Thank you! But I don't understand what you did at the end

Directrix (directrix):

@Genuine --> Do you follow how I got the radius of the circle?

OpenStudy (anonymous):

Yes!

Directrix (directrix):

When you find arc length, what formula do you use?

OpenStudy (anonymous):

θ/360 x circumference

OpenStudy (anonymous):

@Directrix: But what does AL mean?

Directrix (directrix):

AL is arc length as in Arc Length = θ/360 x circumference AL = 30/360 * 2 pi (r) = AL = 1/12 * 6.28 ( 12.36) = 6.47

OpenStudy (anonymous):

OH! Now i understand

OpenStudy (anonymous):

@Directrix: I see what I had done wrong before, instead of square rooting 480 to find the radius I just divided the equation... Thank you so much for helping me (: